Mountain Dental Services is a specialized dental practice whose only service is filling cavities. Mountain has recorded the following for the past nine months:

Month Number of Cavities Filled Total Cost
January 550 $6,050
February 375 5,200
March 325 5,000
April 500 5,750
May 400 5,400
June 650 6,400
July 625 6,200
August 600 6,200
September 300 5,000


Required:
1.
Use the high-low method to estimate total fixed cost and variable cost per cavity filled. (Round your answers to 2 decimal places.)



2. Using these estimates, calculate Mountain’s total cost for filling 450 cavities.
Answer #1

Variable Cost = (Cost at highest activity - Cost at lowest activity) /(Highest activity-Lowest activity)

= (6400-5000)/(650-300)

= 4

Fixed cost = 6400 - (650*4) = 3800

Fixed cost 3800
Variable cost 4

.

Estimated cost = 3800 + (4*450)

= 5600